OpenStudy (anonymous):

There are 4 different routes a student can bike from his house to school. In how many ways can he make a round trip if he uses a different route coming than going 4! ?

OpenStudy (anonymous):

12

OpenStudy (anonymous):

1st2nd,1st3rd,1st4th, 2nd1st,2nd3rd,2nd4th 3rd1st,3rd2nd,3rd4th 4th1st,4th2nd,4th3rd

OpenStudy (anonymous):

4P2=4!/(4-2)! = 4*3 =12

OpenStudy (anonymous):

raheen,good explain ty